# Why They Link: An Intent Taxonomy for Including Hyperlinks in Social Posts

### v1

@@ Abstract.
URLs serve as bridges between social media platforms and the broader web, linking user-generated content to external information resources. On Twitter (X), approximately one in five tweets contains at least one URL, underscoring their central role in information dissemination. While prior studies have examined the motivations of authors who share URLs, such author-centered intentions are difficult to observe in practice. To enable broader downstream use, this work investigates reader-centered interpretations—how users perceive the intentions behind hyperlinks included in posts.We develop an intent taxonomy for including hyperlinks in social posts through a hybrid approach that begins with a bottom-up, data-driven process using large-scale crowdsourced annotations, and is then refined using large language model (LLM) assistance to generate descriptive category names and precise definitions. The final taxonomy comprises 6 top-level categories and 26 fine-grained intention classes, capturing diverse communicative purposes. Applying this taxonomy, we annotate and analyze 1,000 user posts, revealing that advertising, arguing, and sharing are the most prevalent intentions. We further compare our taxonomy with existing taxonomies and demonstrate its utility in a microblog retrieval task by incorporating intent as an additional feature. Overall, our taxonomy provides a foundation for intent-aware information retrieval and NLP applications, enabling more accurate retrieval, recommendation, and interpretation of social media content.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Query intent
@@ 1. Introduction
Communication scientists have developed several research agendas seeking to address the long-standing question “What do people do with media?” (Katz, 1959). In traditional mass communication (e.g., television, radio, newspapers), this inquiry gave rise to the uses and gratifications research tradition (Ruggiero, 2000), an audience-centered framework explaining how individuals engage with media to satisfy specific needs such as information seeking, entertainment, or social connection (Holbert, 2014; Haridakis and Humphries, 2019). From an Information Retrieval (IR) perspective, these needs can be interpreted as information-seeking behaviors, where media use serves as a way of discovering, consuming, or disseminating information objects. With the rise of social media platforms, users are no longer only consumers of content but also content creators (Chen, 2011; Phua et al., 2017; Smock et al., 2011). This shift implies that users simultaneously act as query issuers and document providers.
@@
Social media posts contain not only textual content but also structural and semantic elements such as hashtags, emojis, images, and hyperlinks (or URLs) (Yang et al., 2012; Vempala and Preotiuc-Pietro, 2019; Hu et al., 2017). Existing research has examined social media content at different levels, including general tweet intent (Java et al., 2007), topic-focused messages (Mohammad et al., 2015), and component-specific studies such as hashtags or emojis (Hu et al., 2017). It has been shown that hashtags, emojis, and images can function as rich indexing and query-expansion features in social media search (aka, microblog retrieval) (Wang et al., 2023; Basu et al., 2018) and recommendation tasks (Ding et al., 2013; Hong et al., 2024; Ma and Fan, 2022). However, URLs, which directly link social media posts to external web documents, have received comparatively less attention. Yet, URLs carry particularly strong retrieval implications: they act as pointers to external information resources and encode users’ intentions for exposing or consuming content. Understanding why users include URLs is important for a range of IR tasks, from intent-aware microblog retrieval ranking (Cambazoglu et al., 2021; Li et al., 2023) to credibility assessment and crisis informatics applications (Purohit et al., 2014; Tanaka et al., 2013; Horawalavithana et al., 2021).
@@
Our goal is therefore to construct a comprehensive, empirically grounded taxonomy of user intentions behind URL sharing on microblog platforms, such as Twitter (X). Prior studies have primarily explored this question from the authors’ perspective, using surveys or interviews to examine their motives for including URLs in posts (Holton et al., 2014; Baek et al., 2011). While such approaches capture authentic creator intentions, they are difficult to generalize and impractical for large-scale application, as authors’ true motives are rarely observable in real-world settings. We focus instead on intentions from the perspective of readers—that is, how other users interpret the purpose of including a URL. Inferance of URL intention is not straightforward. The meaning of a URL often cannot be deduced without considering both the post’s text and the linked content. For example, a post that links to an Amazon product web page may appear promotional if we ignore the post’s text (Figure 1). However, the text of the post reveals that the intention is not to advertise the product but rather to solicit the readers’ feedback about it. In another case, a post may claim to provide free streaming for a sporting event, yet the URL leads only to a news site, exposing the latent intention as clickbait (Figure 2). The latter is an example promoting affiliated websites (Cable and Mottershead, 2018; Chakraborty et al., 2017). These examples highlight the need for an intent taxonomy for hyperlinked posts. A well-defined taxonomy can provide a foundation for intent-aware indexing, ranking, recommendation, and credibility assessment (Alexander et al., 2022; Cambazoglu et al., 2021; Pasi and Viviani, 2020). By distinguishing between intentions such as information sharing, promotion, discussion, or deception, IR systems can better model relevance, improve user experience, and support downstream applications such as credibility detection and crisis response (Imran et al., 2014; Wang et al., 2024).
@@
Figure 2. An example of a tweet that promises a free streaming for a sport match throught the URL.
@@ 2.1. Scope and Gaps in Existing Taxonomies
2.2. The Utility of a URL Intentions Framework
@@
Crisis Response In the domain of crisis response and emergency awareness, extensive research has been conducted that leverages posts and other forms of social media content (Pipek et al., 2014). Several studies pay attention to the presence of URLs in crisis-related tweets, with some reporting that over 50% of the tweets in their datasets contain URLs (Purohit et al., 2014; Hughes and Palen, 2009; Meesters et al., 2016; Hunt et al., 2020). For IR systems that support crisis informatics, distinguishing URL intentions is critical. A taxonomy enables retrieval pipelines to prioritize actionable links (e.g., shelter information, emergency hotlines) while downranking malicious or irrelevant ones. It also aids in credibility assessment by signaling whether URLs are shared to inform, to seek help, or to mislead (Horawalavithana et al., 2021). Such intent-aware retrieval can improve the timeliness and accuracy of information accessible to crisis responders.
Misinformation The proposed intent taxonomy can serve as a key resource in advancing misinformation mitigation. By systematically categorizing the intentions underlying URL-sharing, it provides a principled framework for detecting and characterizing malicious or deceptive content. This facilitates both retrospective analyses of misinformation propagation and the design of proactive interventions grounded in credible information dissemination. Thus, the taxonomy helps promote credible information, contributing to a healthier, more informed digital ecosystem.
NLP and IR Tasks In many NLP tasks, such as sentiment analysis, URLs are often treated as peripheral and removed from tweets without exploiting their potential signals (Chong et al., 2014; Shivakumar, 2021). We argue that a taxonomy of URL intentions enables these signals to be reintroduced as interpretable features for both NLP and IR applications. The insights derived from URL intentions offer signals of the communicative purpose, quality, and utility of a post. For instance, distinguishing whether a URL is included to promote, to inform, or merely to reference can enhance sentiment analysis (Jiang et al., 2011), stance detection (Zubiaga et al., 2016), and user modeling (Tsagkias and Blanco, 2012). From an IR perspective, intent-aware representations support richer query understanding, improve document (microblog) ranking, and lead to more precise retrieval (Cambazoglu et al., 2021; Li et al., 2023). As our demonstration reports (Section 5.2), recognizing intention allows retrieval systems to better estimate which posts are valuable for a given query, thereby increasing confidence in their relevance.
@@ 3. Developing the Intent Taxonomy
The large fraction of social media posts containing URLs (Aljebreen et al., 2021; Suh et al., 2010) highlights the diversity of content being circulated and suggests a wide range of underlying intentions behind URL sharing. We describe our approach for constructing a tailored taxonomy that captures a comprehensive set of perceived intentions behind URL inclusion in posts in this section.
@@ 3.1.1. Dataset
Our Tweets with URLs dataset consists of a random collection of tweets. We deliberately avoid using existing tweet datasets, as many of them suffer from one of two limitations: they either remove URLs during normalization or retain URLs that are no longer accessible, often redirecting to error pages or generic homepages. To address these issues, we collect recent tweets and simultaneously crawl the linked web documents, increasing the likelihood of preserving valid tweet–URL pairs. Access to the linked content is essential for our study, as annotators are instructed to consult the URL content when the tweet text alone is insufficient to infer intent. The Tweets with URLs dataset contains over 500,000 tweets. From this collection, we randomly sample 2,500 tweets for taxonomy development.
@@ 3.1.2. Participants
To improve the reliability of crowdsourced annotations, we employed a qualification task to screen AMT workers. The task consisted of annotating five carefully selected tweets and was restricted to AMT Master workers. Workers were evaluated on their understanding of the task and their ability to identify the intended purpose of URL sharing. Of the 200 workers who completed the qualification task, 143 met our acceptance criterion by correctly annotating at least four tweets marked as eligible. Only eligible workers were allowed to participate in the subsequent taxonomy development.
@@ 3.2. Procedures and Results
(A) Collecting. We began our study with a crowdsourcing task on AMT, where participants labeled tweets with their perceived intentions behind URL inclusion. Crowdsourcing provides a scalable and cost-effective solution for gathering a large volume of data from diverse individuals (Cocos et al., 2017; Bloodgood and Callison-Burch, 2010). We opted for an open-coding approach, allowing AMT workers to provide codes or intention classes for tweets using an open-text form. Open coding is a popular method in qualitative research and ground theory which has been used in several works to construct taxonomies in other problems (Scott et al., 2012). This method allowed us to capture a wide range of potential intention classes from the crowd and ensured that workers had the flexibility to express their thoughts and observations without being restricted by predefined categories.
@@
(B) Refining. We refine 2,500 intention classes in the previous step into a final list of 754 intention codes. The process is shown as follows:
@@
Purging: We performed a thorough filtering process to ensure the quality and reliability of the results. We specifically looked for unclear codes or ones that do not appear as actual intentions codes but rather appear as descriptive statements or explanations. For example, the code ”talk and complain about rent increase” is a statement about the tweet itself rather than an intention code about the URL, which in that case should have been “to support an argument.” Out of the 2,500 annotations in the list, we deemed 30 instances unsuitable and discarded those codes. We focus on the annotations that provided clear and concise intentions codes.
@@
We used a simple majority to process the final categories and other decisions. Based on the majority votes, 311 were discarded given their similarity to other codes. In the grouping results, some responses formed more categories than others. The number of groups ranged from 17 to 31 in the responses. We compiled those groups into 28 intentions groups with a total of 442 fine-grained intentions. After the initial grouping, we attempted to establish higher-level classes of intentions. To achieve this, we conducted another round of grouping with a new set of 5 AMT workers. As a result, we generated a shorter list of 6 coarse-grained intention classes (Table 1).
@@ 3.3. A Hybrid Construction Method
We present our intent taxonomy for including hyperlinks in social posts [^1] The complete taxonomy with representative examples will be released after the paper is accepted. in Table 1. The final taxonomy represents a synthesis of human annotation and LLM-based refinement. Specifically, we enhanced the taxonomy derived from crowdsourced annotations by prompting GPT-5 to propose more descriptive category names and clearer definitions, using the crowdsourced category categories and fine-grained intent classes as input. The LLM-generated suggestions were then reviewed and consolidated through a consensus discussion among the research team to ensure interpretability, consistency, and empirical grounding. We also sought to enrich each fine-grained intention class with both illustrative examples and example posts. Prior studies and our experience indicate that annotators prefer taxonomies that include examples alongside definitions when applying them to new annotation tasks (Hettiachchi et al., 2022; Pradhan et al., 2021). Hence, we used LLMs to generate illustrative examples and revised them following the steps described above (see the Illustrative Example column in Table 1). We additionally explored using LLMs to create example posts for each intention class. However, this revealed a key limitation: while LLMs can generate plausible and often representative posts, the examples are synthetic rather than drawn from real-world data. Consequently, all examples in our final taxonomy were manually retrieved and curated from authentic social media posts for each of the 26 intention classes.
@@ 4.1. Dataset
We randomly and uniformly sampled 1,000 tweets from our Tweets with URLs dataset to reduce potential temporal bias and improve the robustness of our taxonomy. Table 2 summarizes key tweet characteristics, including type, length, and total reactions. Tweet type includes replies, quoted tweets, and other types of tweets. For the reactions attribute, we quantify tweet popularity using three engagement signals: likes, replies, and retweets. All metrics were collected more than 24 hours after publication and are assumed to reflect stable engagement levels (Pfeffer et al., 2023).

### v2

@@ Abstract.
URLs serve as bridges between social media platforms and the broader web, linking user-generated content to external information resources. On Twitter (X), approximately one in five tweets contains at least one URL, underscoring their central role in information dissemination. While prior studies have examined the motivations of authors who share URLs, such author-centered intentions are difficult to observe in practice. To enable broader downstream use, this work investigates reader-centered interpretations—how users perceive the intentions behind hyperlinks included in posts. We develop an intent taxonomy for including hyperlinks in social posts through a hybrid approach that begins with a bottom-up, data-driven process using large-scale crowdsourced annotations, and is then refined using large language model (LLM) assistance to generate descriptive category names and precise definitions. The final taxonomy [^1] Our complete intent taxonomy is publicly released at https://github.com/lanfangping/intent-taxonomy-for-hyperlinked-posts comprises 6 top-level categories and 26 fine-grained intention classes, capturing diverse communicative purposes. Applying this taxonomy, we annotate and analyze 1,000 user posts, revealing that advertising, arguing, and sharing are the most prevalent intentions. We further compare our taxonomy with existing taxonomies and demonstrate its utility in a microblog retrieval task by incorporating intent as an additional feature. Overall, our taxonomy provides a foundation for intent-aware information retrieval and NLP applications, enabling more accurate retrieval, recommendation, and interpretation of social media content.
@@
journalyear: 2026copyright: ccconference: 18th ACM Web Science Conference; May 26–29, 2026; Braunschweig, Germanybooktitle: 18th ACM Web Science Conference (WebSci ’26), May 26–29, 2026, Braunschweig, Germanydoi: 10.1145/3795766.3799753isbn: 979-8-4007-2504-3/2026/05ccs: Information systems Blogsccs: Information systems Crowdsourcingccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Communication scientists have developed several research agendas seeking to address the long-standing question “What do people do with media?” (Katz, 1959). In traditional mass communication (e.g., television, radio, newspapers), this inquiry gave rise to the uses and gratifications research tradition (Ruggiero, 2000), an audience-centered framework explaining how individuals engage with media to satisfy specific needs such as information seeking, entertainment, or social connection (Holbert, 2014; Haridakis and Humphries, 2019; He et al., 2020). From an Information Retrieval (IR) perspective, these needs can be interpreted as information-seeking behaviors, where media use serves as a way of discovering, consuming, or disseminating information objects. With the rise of social media platforms, users are no longer only consumers of content but also content creators (Chen, 2011; Phua et al., 2017; Smock et al., 2011). This shift implies that users simultaneously act as query issuers and document providers.
@@
Social media posts contain not only textual content but also structural and semantic elements such as hashtags, emojis, images, and hyperlinks (or URLs) (Yang et al., 2012; Vempala and Preotiuc-Pietro, 2019; Hu et al., 2017). Existing research has examined social media content at different levels, including general tweet intent (Java et al., 2007), topic-focused messages (Mohammad et al., 2015), and component-specific studies such as hashtags or emojis (Hu et al., 2017). It has been shown that hashtags, emojis, and images can function as rich indexing and query-expansion features in social media search (aka, microblog retrieval) (Wang et al., 2023; Basu et al., 2018) and recommendation tasks (Ding et al., 2013; Hong et al., 2024; Ma and Fan, 2022). However, URLs, which directly link social media posts to external web documents, have received comparatively less attention. Our prior work established that tweets containing URLs form a distinct and computationally important class of posts requiring dedicated modeling (Aljebreen et al., 2021). Yet, URLs carry particularly strong retrieval implications: they act as pointers to external information resources and encode users’ intentions for exposing or consuming content. Understanding why users include URLs is important for a range of IR tasks, from intent-aware microblog retrieval ranking (Cambazoglu et al., 2021; Li et al., 2023) to credibility assessment and crisis informatics applications (Purohit et al., 2014; Tanaka et al., 2013; Horawalavithana et al., 2021).
@@
Our goal is therefore to construct a comprehensive, empirically grounded taxonomy of user intentions behind URL sharing on microblog platforms, such as Twitter (X). Prior studies have primarily explored this question from the authors’ perspective, using surveys or interviews to examine their motives for including URLs in posts (Holton et al., 2014; Baek et al., 2011). While such approaches capture authentic creator intentions, they are difficult to generalize and impractical for large-scale application, as authors’ true motives are rarely observable in real-world settings. We instead focus on intentions from the perspective of readers—that is, how other users interpret the purpose of including a URL. Inference of URL-sharing intention is not straightforward. The meaning of a URL often cannot be deduced without considering both the post’s text and the linked content. For example, a post that links to an Amazon product web page may appear promotional if we ignore the post’s text (Figure 1). However, the text of the post reveals that the intention is not to advertise the product but rather to solicit the readers’ feedback about it. In another case, a post may claim to provide free streaming for a sporting event, yet the URL leads only to a news site, exposing the latent intention as clickbait (Figure 2). The latter is an example promoting affiliated websites (Cable and Mottershead, 2018; Chakraborty et al., 2017). These examples highlight the need for an intent taxonomy for hyperlinked posts. A well-defined taxonomy can provide a foundation for intent-aware indexing, ranking, recommendation, and credibility assessment (Alexander et al., 2022; Cambazoglu et al., 2021; Pasi and Viviani, 2020). By distinguishing between intentions such as information sharing, promotion, discussion, or deception, IR systems can better model relevance, improve user experience, and support downstream applications such as credibility detection and crisis response (Imran et al., 2014; Wang et al., 2024).
@@
Figure 2. An example of a tweet that promises a free streaming for a sport match through the URL.
@@ 2.1. Scope and Gaps in Existing Taxonomies
2.2. The Utility of a URL-sharing Intention Framework
@@
Crisis Response In the domain of crisis response and emergency awareness, extensive research has been conducted that leverages posts and other forms of social media content (Pipek et al., 2014). Several studies pay attention to the presence of URLs in crisis-related tweets, with some reporting that over 50% of the tweets in their datasets contain URLs (Purohit et al., 2014; Hughes and Palen, 2009; Meesters et al., 2016; Hunt et al., 2020). For IR systems that support crisis informatics, distinguishing URL-sharing intentions is critical. A taxonomy enables retrieval pipelines to prioritize actionable links (e.g., shelter information, emergency hotlines) while downranking malicious or irrelevant ones. It also aids in credibility assessment by signaling whether URLs are shared to inform, to seek help, or to mislead (Horawalavithana et al., 2021). Such intent-aware retrieval can improve the timeliness and accuracy of information accessible to crisis responders.
Misinformation Our prior work demonstrates that aggregating social media posts by shared URLs reveals structural regularities and dissemination risks that remain invisible at the post level, enabling highly accurate detection of deceptive link-sharing patterns (Aljebreen et al., 2024). Building on this insight, the proposed URL intent taxonomy provides a principled framework for systematically characterizing the motivations behind URL-sharing behaviors. By categorizing the intentions underlying link dissemination, it supports more precise detection of malicious or deceptive campaigns, facilitates retrospective analyses of misinformation propagation, and enables proactive interventions grounded in credible information dissemination. Thus, the taxonomy contributes to promoting trustworthy content and fostering a healthier digital information ecosystem.
NLP and IR Tasks In many NLP tasks, such as sentiment analysis, URLs are often treated as peripheral and removed from tweets without exploiting their potential signals (Chong et al., 2014; Shivakumar, 2021). We argue that a taxonomy of URL-sharing intentions enables these signals to be reintroduced as interpretable features for both NLP and IR applications. The insights derived from URL-sharing intentions offer signals of the communicative purpose, quality, and utility of a post. For instance, distinguishing whether a URL is included to promote, to inform, or merely to reference can enhance sentiment analysis (Wang and Dragut, 2024; Dragut et al., 2010), stance detection (Hosseinia et al., 2019), and user modeling (Tsagkias and Blanco, 2012). From an IR perspective, intent-aware representations support richer query understanding, improve document (microblog) ranking, and lead to more precise retrieval (Cambazoglu et al., 2021; Li et al., 2023; Zhang et al., 2020). As our demonstration reports (Section 5.2), recognizing intention allows retrieval systems to better estimate which posts are valuable for a given query, thereby increasing confidence in their relevance.
@@ 3. Developing the Intent Taxonomy
The large fraction of social media posts containing URLs (Aljebreen et al., 2021; Suh et al., 2010) highlights the diversity of content being circulated and suggests a wide range of underlying intentions behind URL sharing. We describe our approach for constructing a tailored taxonomy that captures a comprehensive set of perceived intentions behind URL inclusion in posts in this section. Further details are available in the supplementary material.
@@ 3.1.1. Dataset
Our Tweets with URLs dataset consists of a random collection of tweets, which was collected using our previously designing API-driven and focused crawling pipelines for large-scale social media and web research (Chen et al., 2024, 2025; Aljebreen et al., 2024, 2021). We deliberately avoid using existing tweet datasets, as many of them suffer from one of two limitations: they either remove URLs during normalization or retain URLs that are no longer accessible, often redirecting to error pages or generic homepages. To address these issues, we collect recent tweets and simultaneously crawl the linked web documents, increasing the likelihood of preserving valid tweet–URL pairs. Access to the linked content is essential for our study, as annotators are instructed to consult the URL content when the tweet text alone is insufficient to infer intent. The Tweets with URLs dataset contains over 500,000 tweets. From this collection, we randomly sample 2,500 tweets for taxonomy development.
@@ 3.1.2. Participants
Given documented concerns about the decline in data quality on Amazon Mechanical Turk (AMT) in recent years— particularly the rise of low-quality and low-fluency responses (Marshall et al., 2023)— we adopted a stricter worker screening process to improve the reliability of crowdsourced annotations. We employed a qualification task to screen AMT workers. The task consisted of annotating five carefully selected tweets and was restricted to AMT Master workers. Workers were evaluated on their understanding of the task and their ability to identify the intended purpose of URL sharing. Out of the 200 workers who completed the qualification task, 143 met our acceptance criterion by correctly annotating at least four tweets and were marked as eligible. Only eligible workers were allowed to participate in the subsequent taxonomy development.
@@ 3.2. Procedures and Results
(A) Collecting. We began our study with a crowdsourcing task on AMT, where participants labeled tweets with their perceived intentions behind URL inclusion. Crowdsourcing provides a scalable and cost-effective solution for gathering a large volume of data from diverse individuals (Cocos et al., 2017; Bloodgood and Callison-Burch, 2010). We opted for an open-coding approach, allowing AMT workers to provide codes or intention classes for tweets using an open-text form. Open coding is a popular method in qualitative research and grounded theory which has been used in several works to construct taxonomies in other problems (Scott et al., 2012). This method allowed us to capture a wide range of potential intention classes from the crowd and ensured that workers had the flexibility to express their thoughts and observations without being restricted by predefined categories.
@@
(B) Refining. We refined 2,500 intention classes in the previous step into a final list of 754 intention codes. The process is shown as follows:
@@
Purging: We performed a thorough filtering process to ensure the quality and reliability of the results. We specifically looked for unclear codes or ones that do not appear as actual intention codes but rather appear as descriptive statements or explanations. For example, the code ”talk and complain about rent increase” is a statement about the tweet itself rather than an intention code about the URL, which in that case should have been “to support an argument.” Out of the 2,500 annotations in the list, we deemed 30 instances unsuitable and discarded those codes. We focus on the annotations that provided clear and concise intention codes.
@@
We used a simple majority to process the final categories and other decisions. Based on the majority votes, 311 were discarded given their similarity to other codes. In the grouping results, some responses formed more categories than others. The number of groups ranged from 17 to 31 in the responses. We compiled those groups into 28 intention groups with a total of 442 fine-grained intentions. After the initial grouping, we attempted to establish higher-level classes of intentions. To achieve this, we conducted another round of grouping with a new set of 5 AMT workers. As a result, we generated a shorter list of 6 coarse-grained intention classes (Table 1).
@@ 3.3. A Hybrid Construction Method
We present our intent taxonomy for including hyperlinks in social posts [^2] The complete taxonomy with representative examples can be found at the shared link. in Table 1. The final taxonomy represents a synthesis of human annotation and LLM-based refinement. Specifically, we enhanced the taxonomy derived from crowdsourced annotations by prompting GPT-5 to propose more descriptive category names and clearer definitions, using the crowdsourced categories and fine-grained intent classes as input. The LLM-generated suggestions were then reviewed and consolidated through a consensus discussion among the research team to ensure interpretability, consistency, and empirical grounding. We also sought to enrich each fine-grained intention class with both illustrative examples and example posts. Prior studies and our experience indicate that annotators prefer taxonomies that include examples alongside definitions when applying them to new annotation tasks (Hettiachchi et al., 2022; Pradhan et al., 2021). Hence, we used LLMs to generate illustrative examples and revised them following the steps described above (see the Illustrative Example column in Table 1). We additionally explored using LLMs to create example posts for each intention class. However, this revealed a key limitation: while LLMs can generate plausible and often representative posts, the examples are synthetic rather than drawn from real-world data. Consequently, all examples in our final taxonomy were manually retrieved and curated from authentic social media posts for each of the 26 intention classes.
@@ 4.1. Dataset
We randomly and uniformly sampled 1,000 tweets from our Tweets with URLs dataset to reduce potential temporal bias and improve the robustness of our taxonomy. Table 2 summarizes key tweet characteristics, including type, length, and total reactions. Tweet type includes replies, quoted tweets, and other types of tweets. For the reactions attribute, we quantify tweet popularity using three engagement signals: likes, replies, and retweets. All metrics were collected more than 24 hours after publication and are assumed to reflect stable engagement levels (Pfeffer et al., 2023; He et al., 2021).